OpenStudy (anonymous):

A transit company carries about 80 000 riders per day for a fare of $1.25. To obtain more revenue, the management plans to increase the fare. It estimates that for every 5 cent increases in fare, it will lose 1000 riders. What fare would result in the greatest revenue and what would the maximum revenue be?

OpenStudy (anonymous):

Let x be the times the company increases its far price. So lets say they increase by 15 cents than x=3 What is the revenue of the company? The number of riders*fare price. Currently it is: 80 000*1.25 After increasing the price by x*0.05 cents: The riders will decrease by 1000*x and the fare will increase by x*0.05 Thus the new revenue is: (80 000-1000x)*(1.25+x0.05) Expanding the bracket 100 000+4000x-1250x-50x^2 -50x^2 +2750x + 100 000 Take out the common factor (-50 so we will have x^2) -50*(x^2 -55x - 2000) As we are looking for max revenue depending on x we can ignore 50 Max of: - (x^2 -55x - 2000) That is the Min of: x^2 -55x - 2000 Hint (x-27.5)^2=x^2-55x^2 +756.25 x^2 -55x - 2000=(x-27.5)^2 -756.25 -2000 (x-27.5)^2 - 2756.25 Has minumum at x-27.5=0 x=27.5 Lets check 27.5*0.05=1.375 new price is: 1.25+1.375=2.625 80000-27.5*1000= 52500 New revenue: 52500*2.625=137 812.5

OpenStudy (anonymous):

thanks! really appreciate your help:D